OpenStudy (anonymous):

Plz help.. how long would it take you to earn $450 in interest on an investment of $2,500 invested at 1.25% compounded continuously?

OpenStudy (anonymous):

2950 = 2,500 e^.0125t e^.0125t = 2950 / 2500 Take ln both sides: .0125 t = ln ( 1.18) => t = ln (1.18) / .0125 = 13

OpenStudy (anonymous):

thanku